Two stones are thrown vertically upward from the ground, one with three times the initial speed of the other. (a) If the faster stone takes 10 s to return on the ground, how long will it take the slower stone to return? (b) If the slower stone reaches a maximum height of H, how high (in terms of H) will the faster stone go? Assume the two stones are freely falling objects. 


1
Expert's answer
2022-06-13T08:22:53-0400

(a)

"t=2\frac{v_0}{g}"

Hence

"t_2=3t_1=3*10\:\rm s=30\: s"

(b)

"H=\frac{v_0^2}{2g}"

Hence

"H_2=9H_1=9H"